How many grams of CuO aree needed to make 1500 mL of N2 at STP


1
Expert's answer
2021-05-18T07:52:45-0400

2NH+ 3CuO= N2+ 3Cu+ 3H2O

1500mL=1.5L

I mol of nitrogen=22.4l

1.5L=(1.5/22.4)=0.067moles

Mole ratio of Copper ii oxide and nitrogen gas=

CuO:N2=3:1=?:0.067

Moles of CuO=3X0.067=0.201

Molar mass of CuO=79.55 grams

moles of CuO used in the reaction=79.55x0.201

15.99 grams